OpenStudy (anonymous):

The base of a square pyramid has a side length of 6 inches. The height of the pyramid is 12 inches. What is the volume of the pyramid? A. 144 in3 B. 72 in3 C. 36 in3 D. 108 in3

OpenStudy (hoblos):

the area of the square base = side^2 once you calculate the area of the base you apply the same previous formula

OpenStudy (anonymous):

it's 72

OpenStudy (anonymous):

Find the area of the square first as : 6^2 and the multiply it with the height.. and the divide the whole above answer by 3 .. So, what do you get @lee_actress

OpenStudy (anonymous):

Is my answer correct?

OpenStudy (hoblos):

can you recalculate it ?

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

No, not :-(

OpenStudy (hoblos):

\[\frac{ 6^2 \times 12 }{ 3 } = ?\]

OpenStudy (anonymous):

144

OpenStudy (anonymous):

@hoblos

OpenStudy (hoblos):

correct

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

any time :)